# Semantically Diverse Language Generation for Uncertainty Estimation in Language Models

### v1

@@ Appendix A Broader Impact
It is important to note that our method evaluates uncertainty based on the information available to the language model. Therefore, it may inaccurately deem a factually incorrect answer as certain if the model’s knowledge base contains similar errors. This issue, often perceived as model “hallucination”, is not a reflection of the model’s uncertainty, but rather a result of factual inaccuracies in the underlying data that is additional to hallucinations due to uncertainty.
@@ Appendix B On the proposal distribution induced by SDLG
In the following, we analyze the proposal distribution induced by SDLG. We consider a probabilistic transformation of one output sequence $\bm{y}^{\prime}$ into another output sequence $\bm{y}$, given by $p(\bm{y}\mid\bm{y}^{\prime},\bm{x},\bm{w})$. This is introduced, because we have to sum over all possible output sequences $\bm{y}^{\prime}$ that we could apply SDLG on, leading to
@@
$p(\bm{y}\mid\bm{y}^{\prime},\bm{x},\bm{w})$ $=\ \sum_{t=1}^{T}\ p(t\mid\bm{y}^{\prime},\bm{x},\bm{w})\ p(\bm{y}\mid t,\bm{y}^{\prime},\bm{x},\bm{w})\ .$ (15)
The construction of $\bm{y}$ from $\bm{y}^{\prime}$ only changes one element of $\bm{y}^{\prime}$ at position $t$ and then generates the postfix new. Therefore, we have $p(\bm{y}\mid t,\bm{y}^{\prime},\bm{x},\bm{w})=0$ for $\bm{y}^{\prime}_{<t}\not=\bm{y}_{<t}$. Consequently, $\sum_{\bm{y}^{\prime}\in\mathcal{Y}}p(\bm{y}^{\prime}\mid\bm{x},\bm{w})$ can be reduced to $\sum_{\bm{y}^{\prime}_{>t}\in\mathcal{Y}_{>t}}p(\bm{y}^{\prime}_{>t}\mid\bm{y}_{\leqslant t},\bm{x},\bm{w})\ p(\bm{y}_{\leqslant t}\mid\bm{x},\bm{w})$ if the factor $p(\bm{y}\mid t,\bm{y}^{\prime},\bm{x},\bm{w})$ is present. There is only one possibility for the prefix with $\bm{y}^{\prime}_{<t}=\bm{y}_{<t}$.
@@
where we used $p(y_{t}\mid\bm{y}_{<t},\bm{x},\bm{w})=1$, since SDLG chooses $y_{t}$ deterministically given $\bm{y}_{<t}=\bm{y}^{\prime}_{<t}$. We assume that all probability mass in $p(t\mid y_{t},\bm{y}_{<t},\bm{x},\bm{w})$ is at the actually observed $t$. This means, given all possible $\bm{y}^{\prime}_{>t}$, $t$ is the most probable position to induce a semantic change. This is a strong assumption, that needs further investigation in future work. Under this assumption, the final result in Eq. (16) reduces to
@@ Appendix C Estimating the Semantic Entropy
As already established in the main paper, directly using the estimator for semantic entropy in Eq. (5) is not possible, because the distribution $p(c\mid\bm{x},\bm{w})$ is not known. Inspecting the implementation of Kuhn et al. [2023] reveals that their estimator of the semantic entropy is using the estimate of the semantic cluster probability distribution
$p(c\mid\bm{x},\bm{w})$ $\approx\ \sum_{n=1}^{N}\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}\ p(\bm{y}^{n}\mid\bm{x},\bm{w})\ ,\quad\bm{y}^{n}\sim p(\bm{y}\mid\bm{x},\bm{w})\ .$ (19)
@@
$\mathbf{\mathrm{H}}(p(c\mid\bm{x},\bm{w}))$ $\approx\ -\frac{1}{M}\sum_{m=1}^{M}\log p(c_{m}\mid\bm{x},\bm{w})\ .$ (20)
@@ C.2 Details on our Semantic Entropy Estimator
$p(c\mid\bm{x},\bm{w})$ $\approx\ \frac{1}{N}\ \sum_{n=1}^{N}\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}\ \frac{p(\bm{y}^{n}\mid\bm{x},\bm{w})}{\hat{q}(\bm{y}^{n}\mid\bm{x},\bm{w})}\ ,\quad\bm{y}^{n}\sim\hat{q}(\bm{y}\mid\bm{x},\bm{w})\ .$ (21)
As this distribution is known by design and can be enumerated, Eq. (21) simplifies to a weighted sum. The quality of this approximator strongly depends on the empirical distribution. Therefore, Eq. (21) should only be used in favor of Eq. (6) if $\{\bm{y}^{n}\}_{n=1}^{N}$ containts output sequences that have very high probability under $p(\bm{y}\mid\bm{x},\bm{w})$. The more these distributions differ, the higher the variance of the estimator, therefore, the lower the approximation quality. We utilized Eq. (21) instead of Eq. (6) for the baseline using multinomial sampling and in addition to the importance sampling we perform with SDLG (c.f. Eq.(8)).
To illustrate the validity of using the estimator in Eq. (21), consider the following example: Given are a probability distribution $p(y)=(0.15,0.1,0.3,0.45)$. Furthermore, the cluster probability distribution $p(c)=(0.25,0.75)$ is derived from this distribution, thus $y_{0},y_{1}\in c_{0}$ and $y_{2},y_{3}\in c_{1}$. The distributions are shown in Fig. 2(a). In Fig. 2(b), we compare the MC approximation using Eq. (6) and Eq. (21). The results show, that the estimator using Eq. (21) is prone to be more biased for a low number of samples, but decreases its variance much faster.
@@ Appendix D Further Experiments and Details
The code and data is available at https://github.com/ml-jku/SDLG.
@@ Results.
A comprehensive overview of all conducted experiments is given in Tab. 2. The results of our reimplementation of methods PE, LN-PE, and SAR match closely with the results reported in prior work [Kuhn et al., 2023, Duan et al., 2023]. Furthermore, we did an in-depth comparison of the two best semantic entropy estimators $\text{SE}^{(*)}_{MS}$ and $\text{SE}^{(*)}_{\texttt{SDLG}{}}$ for all three considered correctness metrics using an extensive range of correctness thresholds, as well as over the number of sampled output sequences. Results on the three considered datasets for all models are depicted in Fig. 3 - 5.
The closest gap between previous methods and SDLG is on the CoQA dataset. Therefore, we more closely inspect the relation between the performance of the methods and the sample size for all methods on this dataset. The results are shown in Fig. 7(a). SDLG shows strong performance already after sampling a single additional output sequence to the initial output sequence $\bm{y}^{\prime}$. Furthermore, inspecting the number of semantic clusters in Fig. 7(b), it can be observed that SDLG finds many more semantic clusters than multinomial sampling.
@@
The computational expenses of SDLG and previous methods is shown in Fig. 6 across number of samples generated by a 2.7 billion parameter language model. It can be observed that the advantage of our method in terms of computational efficiency further increases with an increasing number of samples and longer input and output sequences. To further reduce the computational expenses of our method, we decrease the number of token scores that have to be computed by implementing a token probability threshold of $0.001$, under the rationale that tokens falling below this probability threshold would, in any case, be assigned a very low importance score.
All experiments were performed on four A100 (80GB) GPUs, used independently for small model sizes and together for larger model sizes. With efficient implementations, the experiments took fewer than 1000 GPU hours to execute.
@@ Appendix E Insights into SDLG
Fig. 8 considers the input sequence “Who proposed the theory of relativity?” with the given output sequence “Albert Einstein did.”. When investigating the alternative tokens it becomes clear that not every substitution leads to a change in semantic meaning. It is important to substitute tokens that also receive a high score for altering the semantics. In this example, it is the initial token corresponding to “Einstein” and the alternative token corresponding to “Schweitzer”. Yet, this alone does not directly indicate a high level of uncertainty about the output sequence. High uncertainty should be attributed only if the new output sequence is completed and still has a different semantic meaning. If the language model is uncertain about the originator of the theory of relativity, it completes the new output sequence like “Albert Schweitzer proposed the theory of relativity.”. This would suggest a high uncertainty estimate. However, if the language model is confident about the originator of the theory of relativity, it completes the new output sequence like “Albert Schweitzer didn’t, but Albert Einstein did.”. It is in favor of a low uncertainty estimate since the model reinforces the original semantics. This illustrates that solely considering the predictive uncertainty on a token level is insufficient. Steering the generation towards a different semantics and then continuing the usual generation can be viewed as stress testing the language model.
@@
Fig. 9 shows the computational flow of how the three scores Importance, Substitution, and Attribution are computed for one specific token pair.
@@
We provide examples of output sequences generated by SDLG and standard multinomial sampling on questions from the TruthfulQA dataset in Tab. 3.
Table 2: Comprehensive results: AUROC using different uncertainty measures as a score to distinguish between correct and incorrect answers, using ten samples each. SEMS uses the estimator implemented by Kuhn et al. [2023] while ${}_{\cdots}^{(*)}$ use the proper semantic entropy estimator as introduced in Sec. 3.
$\mathcal{D}$ Metric Model LN-PE PE SAR SEMS SE()${}_{MS}^{(*)}$ SE()${}_{DBS}^{(*)}$ SE()${}_{\texttt{SDLG}{}}^{(*)}$
Correctness Threshold ($\bm{\geq}$) 0.3 0.5 1.0 0.3 0.5 1.0 0.3 0.5 1.0 0.3 0.5 1.0 0.3 0.5 1.0 0.3 0.5 1.0 0.3 0.5 1.0
TruthfulQA OPT-2.7b .438 .439 .463 .523 .516 .538 .604 .611 .629 .408 .405 .416 .786 .860 .865 .647 .686 .708 .855 .920 .934
@@
Figure 6: Computational Efficiency: Average number of Teraflops required across instances for an increasing number of samples generated with SDLG vs. standard multinomial sampling (MS) and Shifting Attention to Relevance (SAR).
Figure 7: CoQA dataset: (a) AUROC using different uncertainty measures across various numbers of samples as a score to distinguish between correct and incorrect answers. Dotted lines indicate the use of the estimator as implemented by Kuhn et al. [2023]. Solid lines indicate the use of the proper estimator as introduced in Sec. 3, if applicable. (b) Average number of semantic clusters found across various numbers of samples. SDLG finds more semantic clusters than multinomial sampling (MS).
@@
Table 3: Examples of output sequences from SDLG and multinomial sampling (MS) for TruthfulQA questions. Bolded segments in answers from SDLG show differences from the initial output sequence.

### v2

@@ Appendix A Broader Impact
It is important to note that our method evaluates uncertainty based on the information available to the language model. Therefore, it may inaccurately deem a factually incorrect answer as certain if the model is consistently trained on such erroneous facts. This issue, often perceived as model “hallucination”, is not a reflection of the model’s uncertainty, but rather a result of factual inaccuracies in the underlying data that is additional to the hallucinations stemming from uncertainty.
@@ Appendix B On the proposal distribution induced by SDLG
In the following, we analyze the proposal distribution induced by SDLG. We consider a probabilistic transformation of one output sequence $\bm{y}^{\prime}$ into another output sequence $\bm{y}$, given by $p(\bm{y}\mid\bm{y}^{\prime},\bm{x},\bm{w})$. This is introduced because we have to sum over all possible output sequences $\bm{y}^{\prime}$ that we could apply SDLG on, leading to
@@
$p(\bm{y}\mid\bm{y}^{\prime},\bm{x},\bm{w})\$$=\ \sum_{t=1}^{T}\ p(t\mid\bm{y}^{\prime},\bm{x},\bm{w})\ p(\bm{y}\mid t,\bm{y}^{\prime},\bm{x},\bm{w})\ .$ (15)
The construction of $\bm{y}$ from $\bm{y}^{\prime}$ only changes one element of $\bm{y}^{\prime}$ at position $t$ and then generates the new postfix. Therefore, we have $p(\bm{y}\mid t,\bm{y}^{\prime},\bm{x},\bm{w})=0$ for $\bm{y}^{\prime}_{<t}\not=\bm{y}_{<t}$. Consequently, $\sum_{\bm{y}^{\prime}\in\mathcal{Y}}p(\bm{y}^{\prime}\mid\bm{x},\bm{w})$ can be reduced to $\sum_{\bm{y}^{\prime}_{>t}\in\mathcal{Y}_{>t}}p(\bm{y}^{\prime}_{>t}\mid\bm{y}_{\leqslant t},\bm{x},\bm{w})\ p(\bm{y}_{\leqslant t}\mid\bm{x},\bm{w})$ if the factor $p(\bm{y}\mid t,\bm{y}^{\prime},\bm{x},\bm{w})$ is present. There is only one possibility for the prefix with $\bm{y}^{\prime}_{<t}=\bm{y}_{<t}$.
@@
where we used $p(y_{t}\mid\bm{y}_{<t},\bm{x},\bm{w})=1$, since SDLG chooses $y_{t}$ deterministically given $\bm{y}_{<t}=\bm{y}^{\prime}_{<t}$. We assume that all probability mass in $p(t\mid y_{t},\bm{y}_{<t},\bm{x},\bm{w})$ is at the actually observed $t$. This means, given all possible $\bm{y}^{\prime}_{>t}$, $t$ is the most probable position to induce a semantic change. This is a strong assumption that needs further investigation in future work. Under this assumption, the final result in Eq. (16) reduces to
@@ Appendix C Estimating the Semantic Entropy
As already established in the main paper, directly using the estimator for semantic entropy in Eq. (5) is not possible because the distribution $p(c\mid\bm{x},\bm{w})$ is not known. Inspecting the implementation of Kuhn et al. [2023] reveals that their estimator of the semantic entropy is using the estimate of the semantic cluster probability distribution
$p(c\mid\bm{x},\bm{w})\$$\approx\ \sum_{n=1}^{N}\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}\ p(\bm{y}^{n}\mid\bm{x},\bm{w})\ ,\quad\bm{y}^{n}\sim p(\bm{y}\mid\bm{x},\bm{w})\ .$ (19)
@@
$\mathbf{\mathrm{H}}(p(c\mid\bm{x},\bm{w}))\$$\approx\ -\frac{1}{M}\sum_{m=1}^{M}\log p(c_{m}\mid\bm{x},\bm{w})\ .$ (20)
@@ C.2 Details on our Semantic Entropy Estimator
$p(c\mid\bm{x},\bm{w})\$$\approx\ \frac{1}{N}\ \sum_{n=1}^{N}\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}\ \frac{p(\bm{y}^{n}\mid\bm{x},\bm{w})}{\hat{q}(\bm{y}^{n}\mid\bm{x},\bm{w})}\ ,\quad\bm{y}^{n}\sim\hat{q}(\bm{y}\mid\bm{x},\bm{w})\ .$ (21)
As this distribution is known by design and can be enumerated, Eq. (21) simplifies to a weighted sum. The quality of this approximator strongly depends on the empirical distribution. Therefore, Eq. (21) should only be used in favor of Eq. (6) if $\{\bm{y}^{n}\}_{n=1}^{N}$ containts output sequences that have very high probability under $p(\bm{y}\mid\bm{x},\bm{w})$. The more these distributions differ, the higher the variance of the estimator, and therefore, the lower the approximation quality. We utilized Eq. (21) instead of Eq. (6) for the baseline using multinomial sampling and in addition to the importance sampling we perform with SDLG (c.f. Eq.(8)).
To illustrate the validity of using the estimator in Eq. (21), consider the following example: Given are a probability distribution $p(y)=(0.15,0.1,0.3,0.45)$. Furthermore, the cluster probability distribution $p(c)=(0.25,0.75)$ is derived from this distribution, thus $y_{0},y_{1}\in c_{0}$ and $y_{2},y_{3}\in c_{1}$. The distributions are shown in Fig. 5(a). In Fig. 5(b), we compare the MC approximation using Eq. (6) and Eq. (21). The results show that the estimator using Eq. (21) is prone to be more biased for a low number of samples but decreases its variance much faster.
@@ Appendix D Further Experiments and Details
The code and data are available at https://github.com/ml-jku/SDLG.
Hyperparameters.
For baseline methods, we performed an extensive hyperparameter search for each dataset with the SEMS temperature $\in\{0.25,0.5,1.0,1.5,2.0\}$ and the SEDBS penalty term $\in\{0.2,0.5,1.0\}$. Also, each method uses 10 generations to assign an uncertainty estimator, as prior work suggests that sample sizes above 10 do not significantly improve the performance of the uncertainty measures [Kuhn et al., 2023, Duan et al., 2023].
@@ Results.
A comprehensive overview of all conducted experiments is given in Tab. 2. The results of our reimplementation of methods PE, LN-PE, and SAR match closely with the results reported in prior work [Kuhn et al., 2023, Duan et al., 2023]. Furthermore, we did an in-depth comparison of the two best semantic entropy estimators $\text{SE}^{(*)}_{MS}$ and $\text{SE}^{(*)}_{{\texttt{SDLG}}{}}$ for all three considered correctness metrics using an extensive range of correctness thresholds, as well as over the number of sampled output sequences. Results on the three considered datasets for all models are depicted in Fig. 6 - 8.
@@
The computational expenses of SDLG and previous methods are shown in Fig. 4 across a number of samples generated by a 2.7 billion parameter language model. It can be observed that the advantage of our method in terms of computational efficiency further increases with an increasing number of samples and longer input and output sequences. To further reduce the computational expenses of our method, we decrease the number of token scores that have to be computed by implementing a token probability threshold of $0.001$, under the rationale that tokens falling below this probability threshold would, in any case, be assigned a very low importance score.
@@ Appendix E Insights into SDLG
Fig. 10 considers the input sequence “Who proposed the theory of relativity?” with the given output sequence “Albert Einstein did”. When investigating the alternative tokens it becomes clear that not every substitution leads to a change in semantic meaning. It is important to substitute tokens that also receive a high score for altering the semantics. In this example, it is the initial token corresponding to “Einstein” and the alternative token corresponding to “Schweitzer”. Yet, this alone does not directly indicate a high level of uncertainty about the output sequence. High uncertainty should be attributed only if the new output sequence is completed and still has a different semantic meaning. If the language model is uncertain about the originator of the theory of relativity, it completes the new output sequence like “Albert Schweitzer proposed the theory of relativity”. This would suggest a high uncertainty estimate. However, if the language model is confident about the originator of the theory of relativity, it completes the new output sequence like “Albert Schweitzer didn’t, but Albert Einstein did”. It is in favor of a low uncertainty estimate since the model reinforces the original semantics. This illustrates that solely considering the predictive uncertainty on a token level is insufficient. Steering the generation towards a different semantics and then continuing the usual generation can be viewed as stress-testing the language model.
@@
Fig. 11 shows the computational flow of how attribution, substitution, and importance scores are computed for one specific token pair, a present token (black) as the fourth token of the output sequence and a potential alternative token (turquoise). Initially, the language model takes in the input sequence embeddings (purple) and generates the output sequence by selecting next tokens based on their token probabilities (grey). The token embeddings of the selected tokens (blue) are matched with the embedding space of the NLI model through a Bilingual Mapping, which has to be applied to align the token embeddings of the language model and the NLI model in case the tokenizers are not identical (see Sec. D). The NLI model then predicts that the output sequence entails itself. Given this prediction, a loss is computed to the target “contradiction” so that the gradients of this loss with respect to the token embeddings indicate which part of the output sequence to change in order to get a contradiction, namely a semantically different output sequence. The attribution and substitution scores are based on these gradient vectors (orange), while the importance score is based on the likelihood of the alternative token (turquoise).
@@
Tab. 3 visualizes five randomly selected TruthfulQA questions and their receptive output sequences generated by SDLG and standard multinomial sampling (MS). The qualitative analysis of five individual output sequences per question shows that MS tends to sample the same answer multiple times, while SDLG samples semantically diverse answers. For the first question, the token associated with “July” appears to have a significantly higher probability than the tokens associated with “August” or “September”. As a result, MS predominantly selects “July”. In contrast, SDLG considers not only token likelihoods but also semantic diversity on a sequence level, which leads it to select “August” and “September” despite their lower token likelihoods.
Impact of the three distinct scores.
Fig. 9 shows the performance of SDLG when considering different combinations of attribution, substitution, and importance score. Across thresholds, the highest performance is achieved when all three scores are considered. It can be observed that the substitution score has a greater positive impact than the attribution score, which aligns with the fact that the output sequences still consist of a relatively small number of present tokens. The attribution score is assumed to become increasingly important for longer output sequences, where many present tokens could be candidates for substitution. Since each score positively contributes to the performance of SDLG and the computational cost of computing the scores is negligible compared to sampling output sequences, it is recommended to include all three scores, even for shorter sequences.
[…]
TruthfulQA
OPT-2.7b .438 .439 .463 .523 .516 .538 .604 .611 .629 .408 .405 .416 .786 .860 .865 .647 .686 .708 .855 .920 .934
@@
Figure 9: Ablation study on the three token scores using the OPT-2.7b model with the TruthfulQA dataset. The included scores are weighted equally. Importance score is assumed to always be included, as it is the only score that considers the individual token probabilities (which usually is the only criteria of other sampling methods, e.g. standard multinomial sampling or beam search).
@@
Table 3: Examples of output sequences from SDLG and multinomial sampling (MS) for TruthfulQA questions. Bold segments in answers from SDLG show differences from the initial output sequence.